# MOCHa: Multi-Objective Reinforcement Mitigating Caption Hallucinations

### v1

@@ 6 Conclusion
We have shown that using a multi-objective reward signal allows for optimizing image captioning models to reduce open-vocabulary hallucinations while preserving caption quality. In addition to superior results on standard captioning and hallucination metrics, our OpenCHAIR benchmark demonstrates the utility of our approach in an open-world setting, while providing a new metric for further work on hallucination-aware image captioning. Moreover, our method and metrics may be applied flexibly to a variety of model sizes and architectures.
Appendix
[…]
@@ Appendix A Interactive Visualization
For additional qualitative results, we refer the reader to the interactive visualization tool provided at https://assafbk.github.io/mocha_vis_tool. We provide image captioning results using BLIP-Large with and without MOCHa for 500 randomly selected test images from MS-COCO [23] and Flickr30K [47]. To visually emphasize the hallucination rate in the predictions, for each model we calculate the NLI contradiction probability [^3] Using the same pretrained NLI model described in the main paper. between the top beam and a ground-truth caption (which is depicted below the image), and report the difference in the contradiction probability between the two models. Samples are ordered via n-gram similarity between the predictions of both models, listing the most different predictions first, allowing for better emphasizing items with evident differences first. This is calculated by considering the top 5 beams of BLIP as reference texts and the top 5 beams of BLIP+MOCHa as candidate sentences; we then compute the average BLEU [27] score between each candidate and all references.
@@ B.1 MOCHa Implementation Details
As discussed in Rennie et al. [30], we reduce variance in gradient estimates by shifting the reward function to have zero mean; we apply this to the reward function before adding the KL penalty. To perform this shifting, we subtract the sample mean of this reward (without KL penalty) among all predictions for a given image in a minibatch.
During each training iteration, we build minibatches by selecting 10 images and then generating 10 predictions per image (hence 100 image-prediction pairs total). Generation settings are detailed in the main paper; in addition, we cap generations to be at most 40 tokens. During MOCHa training, we freeze the image encoder of all models, training the text encoder components alone. For BLIP-Large and BLIP-Base we use gradient clipping of 5, learning rate of 1e-6 and 4 PPO steps in each iteration. BLIP-2 is trained with low rank adapters (LoRA) over the keys and values of the decoder attention layers [12] with a learning rate of 1e-6. GIT-base is trained with a learning rate of 1e-5 with 4 PPO steps and gradient clipping of 5.
@@ B.2 OpenCHAIR Implementation details
Generating Diverse Captions  We start by parsing all objects in MS-COCO’s human-annotated captions by first identifying nouns via syntactic parsing [^5] Using the $en\_core\_web\_md$ pipeline from the SpaCy [11] library.. Then, we then filter these for highly concrete nouns, by using the values recorded by Hessel et al. [9] with threshold 4.5. We used these objects, coupled with their corresponding captions, to prompt an instruction-tuned LLM666meta-llama/Llama-2-70b-chat-hf to rephrase the captions with different objects. We used stochastic sampling with top-p of 0.9 and temperature of 0.6 for this LLM generation. While this stage increases the object diversity, we notice that the output still includes many common objects that have a significant overlap with those in MS-COCO. To overcome this issue, we filter out all captions that do not include rare objects, defining an object as rare if its appearance frequency in the dataset is in the lowest 10th percentile. The remaining captions are used as few-shot examples for a LLM777meta-llama/Llama-2-13b (base, not instruction-tuned) to generate new captions, to further increase diversity. We used 10 few shot example for each generated caption, and text is generated using sampling with temperature 0.8. We generate 2,000 captions from the LLM and feed them as prompts to the text-to-image generation model Stable Diffusion XL [29], which generates a single image for each caption. For image generation, we use 40 sampling steps and guidance scale of 10. We also employ negative prompting using the prompt “unclear, deformed, out of image, disfigured, body out of frame” to encourage generation of clear objects in the output images.
@@
B.3 OpenCHAIR Analysis and Comparison to CHAIR
Figure 9: Object Coverage, CHAIR vs. OpenCHAIR. We display the object coverage of CHAIR (over MS-COCO) and OpenCHAIR, measured as the number of unique objects. In OPENChair, objects are found using the parsing method described in Section B.2. As can be observed, the proposed benchmark has significantly greater coverage of different objects.
In Figure 9, we show the difference in object coverage for CHAIR (over MS-COCO) and OpenCHAIR. As can be seen in the figure, OpenCHAIR covers many more objects than CHAIR. This is also reflected qualitatively, as CHAIR is missing many common object types, including daily objects like shoe and guitar. More examples of objects from the OpenCHAIR dataset include: pearl, tiger, sand, tricycle, corkscrew, toy, charcoal, text, pinecone, grandfather, chocolate, wheelchair, wand. A list of all additional OpenCHAIR objects (those not found in CHAIR) can be found in https://github.com/assafbk/mocha_vis_tool/blob/main/openchair_objects.txt.
We also note an additional limitation of CHAIR related to object coverage, namely that CHAIR uses a fixed list of synonyms to coarsely aggregate different, semantically similar objects. For example, the word person is synonymous to policeman, bride, thief, baby, girl, boy, hunter, etc. A prediction that includes one of these words will be considered correct according to CHAIR. By contrast, OpenCHAIR relies on a LLM to determine whether two objects are identical, letting us for instance to consider the objects tuna, salmon and goldfish as distinct. Conversely, this LLM evaluation is robust to valid synonymous relations between different words, e.g. identifying the word fish as valid when it is predicted in place of one of the more fine-grained fish terms.
B.4 TLC-A Comparison Implementation Details
In order to compare our method to TLC-A [28], we received code from its authors and implemented it in our setup. TLC-A is a decoding-time method applied to auto-regressive captioning models, and in our setting we apply it to different models (e.g. BLIP-Large) than those tested by Petryk et al (e.g. OFA). Of particular note is that TLC-A requires selecting a threshold confidence value, which is used in the decoding phase to re-rank generated beams according to the confidence assigned to COCO object tokens. Petryk et al. recommend calibrating this threshold using the COCO validation set to achieve a precision level of at least 99%; however, in our experiments we find that this value cannot be achieved by the models we consider without sacrificing most of the recall, as illustrated in Figure 10. Therefore, we instead use the COCO validation set to select the best-performing threshold with respect to the CHAIR metric, as shown in Table 4. The selected confidence threshold is 0.33 and it achieves a precision of 98.3% and a recall of 84%.
@@ B.4 TLC-A Comparison Implementation Details
Figure 10: Precision-recall curve for selecting TLC-A threshold. As detailed in  [28], we compute a precision-recall curve over the predicted object confidences. As illustrated above, the 99% precision threshold recommended by Petryk et al. [28] cannot be achieved by BLIP-Large on the COCO Karpathy validation set. Hence, in our setting we must adjust the threshold to find a reasonable balance between precision and recall.
@@
Table 4: Selecting a threshold for TLC-A. We evaluate TLC-A with different thresholds (as described by Petryk et al. [28]) over the COCO caption Karpathy validation set. In the first row we have BLIP without TLC-A. We indicate the selected threshold which achieves the best CHAIR scores overall in bold. B@4, C, CHi, CHs, BSc, $\overline{p}$ denote BLEU-4, CIDEr, CHAIR instance and CHAIR sentence, BERTScore, and NLI $p(\text{contr.})$. ) metrics respectively. P, R are the precision and recall that each threshold (for predicted object confidences) achieves over the validation set.
$\emptyset$
a painting of oranges and a silver pitcher on a table
two giraffes eating leaves from a tree
$-r_{kl}$
a painting of some items
some giraffes in the field
$r$
a painting of a pitcher, oranges, and a candle on a table
a giraffe eating leaves from a tree in a field
88footnotetext: Reference ground truth captions: Painting of oranges, a bowl, candle, and a pitcher (left) and A giraffe grazing on a tree in the wilderness with other wildlife (right).
@@
BLIP-L 0.316 0.270 41.5 138.4 244 2.3 3.5 0.244 0.679
BLIP-L+M 0.295 0.259 41.9 139.6 246 2.1 3.1 0.206 0.682
  $-r_{f}$ 0.305 0.267 43.0 142.3 246 2.8 4.4 0.249 0.691
  $-r_{a}$ 0.303 0.257 41.1 132.9 244 1.5 2.3 0.174 0.66
  $-r_{kl}$ 0.241 0.225 27.6 98.9 179 1.4 1.9 0.135 0.62
  $-ppo$ 0.287 0.256 39.4 127.6 229 2.5 3.76 0.212 0.664
@@
A man sitting on a chair with a large stuffed animal, specifically a teady bear, on his lap
a man sitting on a chair holding a large stuffed animal
Figure 12: LLaVa-RLHF V.S MOCHa. We illustrate that RLHF training does not necessarily solve the hallucination problem of VLM models by showing a generation produced by LLaVa-RLHF [38] compared to BLIP+MOCHa. For both models, we use the prompt “a photography of” for generation. See Table 6 for a quantitative comparison.
@@ Appendix C Additional Results
C.1 OpenCHAIR with Beam Search vs. Sampling
We note that OpenCHAIR can be calculated using either beam search ($OCH_{b}$) or sampling ($OCH_{s}$). While $OCH_{b}$ matches the decoding method for deterministic image captioning, $OCH_{s}$ has the advantage of allowing for statistical tests of significance, as its values across samples can be compared between models using a t-test. We reproduce our results with both $OCH_{b}$ and $OCH_{s}$ values reported in Tables 1 (core results) and 5 (ablation results). As is seen in both tables, these scores are highly correlated and are consistent with our general observations regarding hallucinations. Notably, in Table 1 we see that our model improves both open- and closed-vocabulary hallucination metrics, while TLC-A (a closed-vocabulary method) improves the closed-vocabulary CHAIR metric but fails to improve our open-vocabulary $OCH_{b}$.
We also note that the differences in $OCH_{s}$ values before and after MOCHa-tuning in Table 1 are all statistically significant ($p<0.05$), as measured by Welch’s two-sided t-tests applied to the $OCH$ values for each sample between models, and applying Bonferroni multiple testing correction to adjust significance levels for the four pairs of base and MOCHa-tuned models listed.
C.2 Additional Ablations

### v3

@@ 6 Conclusion
We have shown the significance of operating in an open-vocabulary setting to effectively quantify and mitigate caption hallucinations. These are explicitly measured by our OpenCHAIR benchmark, and our MOCHa framework allows for optimizing captioning models to reduce such hallucinations while preserving caption quality. This reduction is demonstrated on our benchmark and other existing metrics. Our method and benchmark may be applied flexibly to a variety of model sizes and architectures, which we foresee providing a framework for future work on hallucination-aware captioning.
7 Limitations
While OpenCHAIR provides diverse coverage of object types, it does not directly measure non-object hallucinations (e.g. hallucinated attributes or relations between entities), which are also targeted by sequence-level approaches such as our MOCHa optimization. We have focused on objects as a natural extension of the existing closed-vocabulary object hallucination benchmark CHAIR, and due to the fact that extracting and comparing objects from image captions is a relatively well-defined task. Future work may consider extending our OpenCHAIR concept to non-objects, specifically, constructing a robust benchmark for evaluating hallucinations on the attribute-, relation-, predicate-level, or of other types, utilizing elements of our methodology such as open-vocabulary LLM evaluation. Furthermore, we acknowledge that captioning models may show different performance on the synthetic images found in OpenCHAIR relative to natural images, although we have found it to correlate empirically to other hallucinations metrics and human intuition.
We emphasize that our work does not solve the hallucination problem completely, although it presents a significant step towards this goal. Note also that we have focused in this work on the image captioning domain, while modern VLMs are often applied to diverse tasks such as VQA and visual instruction-following for which hallucinations also pose a significant challenge. We hope that our proposed strategy will pave the way for future research on hallucination reduction in all of these domains, in which open-vocabulary approaches also present significant promise.
8 Ethics Statement
This work focuses on measuring and mitigating hallucinations in visual-language models (VLMs). As such it is expected to increase the reliability of VLMs and the ability to measure their performance, which is important when using them in real world systems. This is expected to have a positive impact on the use of VLMs in the society. However, we do recognize that the foundation models used in the OpenCHAIR construction and evaluation pipeline and those used to calculate the MOCHa reward function could propagate biases. We anticipate further research into such biases before relying on our work beyond the research environment.
@@ Appendix A Interactive Visualization
For additional qualitative results, we refer the reader to the interactive visualization tool provided at https://assafbk.github.io/mocha_vis_tool.
We provide image captioning results using BLIP-Large with and without MOCHa for 350 randomly selected test images from MS-COCO Lin et al. (2015) and Flickr30K Young et al. (2014).
To visually emphasize the hallucination rate in the predictions, for each model we calculate the NLI contradiction probability [^1] Using the same pretrained NLI model described in the main paper. between the top beam and a ground-truth caption (which is depicted below the image), and report the difference in the contradiction probability between the two models. Samples are ordered via n-gram similarity between the predictions of both models, listing the most different predictions first, allowing for better emphasizing items with evident differences first. This is calculated by considering the top 5 beams of BLIP as reference texts and the top 5 beams of BLIP+MOCHa as candidate sentences; we then compute the average BLEU Papineni et al. (2002) score between each candidate and all references.
@@ B.1 MOCHa Implementation Details
As discussed in Rennie et al. Rennie et al. (2017), we reduce variance in gradient estimates by shifting the reward function to have zero mean; we apply this to the reward function before adding the KL penalty. To perform this shifting, we subtract the sample mean of this reward (without KL penalty) among all predictions for a given image in a minibatch.
During each training iteration, we build minibatches by selecting 10 images and then generating 10 predictions per image (hence 100 image-prediction pairs total). We use nucleus sampling Holtzman et al. (2019) with $p=0.9$ and temperature $t=1.2$, and we cap generations to be at most 40 tokens. We apply PPO reinforcement with clipping parameter $\epsilon=0.2$. For our reward function, we use coefficients $\alpha=0.5$ and $\beta\in[0.004,0.06]$ (depending on the model optimized).
During MOCHa training, we freeze the image encoder of all models, training the text encoder components alone. For BLIP-Large and BLIP-Base we use gradient clipping of 5, learning rate of 1e-6 and 4 PPO steps in each iteration. BLIP-2 is trained with low rank adapters (LoRA) over the keys and values of the decoder attention layers Hu et al. (2021) with a learning rate of 1e-6. GIT-base is trained with a learning rate of 1e-5 with 4 PPO steps and gradient clipping of 5.
@@ B.2 OpenCHAIR Implementation details
Generating Diverse Captions  We start by parsing all objects in MS-COCO’s human-annotated captions by first identifying nouns via syntactic parsing [^3] Using the $en\_core\_web\_md$ pipeline from the SpaCy Honnibal and Montani (2017) library.. We then filter these for highly concrete nouns, by using the values recorded by Hessel et al. Hessel et al. (2018) with threshold 4.5. We used these objects, coupled with their corresponding captions, to prompt an instruction-tuned LLM444meta-llama/Llama-2-70b-chat-hf (4-bit quant.) to rephrase the captions with different objects. We used stochastic sampling with top-p of 0.9 and temperature of 0.6 for this LLM generation. While this stage increases the object diversity, we notice that the output still includes many common objects that have a significant overlap with those in MS-COCO. To overcome this issue, we filter out all captions that do not include rare objects, defining an object as rare if its appearance frequency in the dataset is in the lowest 10th percentile. The remaining captions are used as few-shot examples for a LLM555meta-llama/Llama-2-13b (base, not instruction-tuned) to generate new captions, to further increase diversity. We used 10 few shot example for each generated caption, and text is generated using sampling with temperature 0.8. We generate 2,000 captions from the LLM and feed them as prompts to the text-to-image generation model Stable Diffusion XL Podell et al. (2023), which generates a single image for each caption. For image generation, we use 40 sampling steps and guidance scale of 10. We also employ negative prompting using the prompt “unclear, deformed, out of image, disfigured, body out of frame" to encourage generation of clear objects in the output images.
@@
Figure 11: Precision-recall curve for selecting TLC-A threshold. As detailed in  Petryk et al. (2023), we compute a precision-recall curve over the predicted object confidences. As illustrated above, the 99% precision threshold recommended by Petryk et al. Petryk et al. (2023) cannot be achieved by BLIP-Large on the COCO Karpathy validation set. Hence, in our setting we must adjust the threshold to find a reasonable balance between precision and recall.
Refer to caption Refer to caption Refer to caption
$\emptyset$ a painting of oranges and a silver pitcher on a table two giraffes eating leaves from a tree
$-r_{kl}$ a painting of some items some giraffes in the field
$r$ a painting of a pitcher, oranges, and a candle on a table a giraffe eating leaves from a tree in a field
55footnotetext: Reference ground truth captions: Painting of oranges, a bowl, candle, and a pitcher (left) and A giraffe grazing on a tree in the wilderness with other wildlife (right).
BLIP2 Pred = ‘E’ Pred = ‘H’
GT = ‘E’ 332 42
GT = ‘H’ 0 54
BLIP-L Pred = ‘E’ Pred = ‘H’
GT = ‘E’ 353 44
GT = ‘H’ 0 31
GIT-B Pred = ‘E’ Pred = ‘H’
GT = ‘E’ 325 36
GT = ‘H’ 1 66
OFA-L Pred = ‘E’ Pred = ‘H’
GT = ‘E’ 336 45
GT = ‘H’ 1 46
Table 3: Human Evaluation of OpenCHAIR Benchmark. The tables illustrate a correlation measurement between OpenCHAIR’s automatic hallucination annotations (Pred) and manual human hallucination annotations (GT). ‘E’, ‘H’ stand for ’object Exists’, ’object Hallucinated’, respectively. BLIP2, BLIP-L, GIT-B and OFA-L stand for BLIP2-2.7b, BLIP-Large, GIT-Base, OFA-Large, all fine-tuned for image-captioning over COCO.
BLIP2 Pred = ‘E’ Pred = ‘H’
GT = ‘E’ 416 3
GT = ‘H’ 4 5
BLIP-L Pred = ‘E’ Pred = ‘H’
GT = ‘E’ 413 2
GT = ‘H’ 4 9
GIT-B Pred = ‘E’ Pred = ‘H’
GT = ‘E’ 412 1
GT = ‘H’ 3 12
OFA-L Pred = ‘E’ Pred = ‘H’
GT = ‘E’ 418 2
GT = ‘H’ 3 5
Table 4: Human Evaluation of CHAIR Benchmark. The tables illustrates a correlation measurement between CHAIR’s automatic hallucination annotations (Pred) and manual human hallucination annotations (GT). ‘E’, ‘H’ stand for ’object Exists’, ’object Hallucinated’, respectively. BLIP2, BLIP-L, GIT-B and OFA-L stand for BLIP2-2.7b, BLIP-Large, GIT-Base, OFA-Large, all fine-tuned for image-captioning over COCO.
B.3 LURE Comparison
To evaluate LURE Zhou et al. (2024) in our setup, we followed the authors’ instructions [^6] https://github.com/YiyangZhou/LURE/blob/main/README.md and applied their pre-trained model (YiyangAiLab/LURE, over MiniGPT-4 with VICUNA-13b) to our predicted captions. Both BLIP-L’s and BLIP-2’s predictions (with beam search decoding, 5 beams) were supplied to LURE’s revisor along with the probabilities of each predicted token for the highest scoring beam. After additional parsing, LURE revised BLIP-L’s and BLIP-2’s sentences, which we then evaluated with various metrics. An example of this procedure is given below:
  • Original caption from BLIP-L, masked by LURE’s algorithm: a woman standing in a kitchen with [IDK]
  • LURE instruction: According to the picture, remove the information that does not exist in the following description:
  • LURE output revision: In the image, a smiling woman is standing in a kitchen with various food items spread out on a table. She seems to be posing and enjoying the scene, perhaps taken during a cooking session. The table is filled with food, including several pieces of sushi placed near the center, along with some pasta on the left side, cucumbers near the right side, and a couple of apples towards the back.
B.4 TLC-A Comparison
In order to compare our method to TLC-A Petryk et al. (2023), we received code from its authors and implemented it in our setup. TLC-A is a decoding-time method applied to auto-regressive captioning models, and in our setting we apply it to different models (e.g. BLIP-Large) than those tested by Petryk et al (e.g. OFA). Of particular note is that TLC-A requires selecting a threshold confidence value, which is used in the decoding phase to re-rank generated beams according to the confidence assigned to COCO object tokens. Petryk et al. recommend calibrating this threshold using the COCO validation set to achieve a precision level of at least 99%; however, in our experiments we find that this value cannot be achieved by the models we consider without sacrificing most of the recall, as illustrated in Figure 11. Therefore, we instead use the COCO validation set to select the best-performing threshold with respect to the CHAIR metric, as shown in Table 5. The selected confidence threshold is 0.33 and it achieves a precision of 98.3% and a recall of 84% over the validation set.
@@ B.4 TLC-A Comparison
Table 5: Selecting a threshold for TLC-A. We evaluate TLC-A with different thresholds (as described by Petryk et al. Petryk et al. (2023)) over the COCO caption Karpathy validation set. In the first row we have BLIP without TLC-A. We indicate the selected threshold which achieves the best CHAIR scores overall in bold. B@4, C, CHi, CHs, BSc, $\overline{p}$ denote BLEU-4, CIDEr, CHAIR instance and CHAIR sentence, BERTScore, and NLI $p(\text{contr.})$. ) metrics respectively. P, R are the precision and recall that each threshold (for predicted object confidences) achieves over the validation set.
@@
BLIP-B 24.8 87.5 2.6 2.8 17.6 0.206 0.557
BLIP-B+M (ours) 26.0 91.3 2.2 2.5 16.4 0.176 0.576
BLIP-L 41.5 138.4 2.3 3.5 19.2 0.244 0.679
BLIP-L+M (ours) 41.9 139.6 2.1 3.1 18.3 0.206 0.682
BLIP2 43.4 144.3 1.7 2.6 17.0 0.207 0.684
BLIP2+M (ours) 44.0 144.3 1.4 2.3 16.6 0.199 0.684
GIT-B 38.7 128.1 4.2 2.9 24.7 0.284 0.656
GIT-B+M (ours) 39.0 128.4 3.9 2.7 22.9 0.221 0.657
Table 6: Quantitative results for state-of-the-art VLM models on the COCO Caption Karpathy test set. +M refers to MOCHa. BSc and $\bar{p}$ denote BERTScore and NLI contradiction probability rewards. B@4, C, CH, OCH, BSc and $\overline{p}$ denote BLEU-4, CIDEr, CHAIR (i for instance, s for sentence), OpenCHAIR, BERTScore, and NLI $p(\text{contr.})$. ) metrics respectively. All results are generated by using their officially provided checkpoints and hyperparameters. Best results are shown in bold.
[…]
BLIP-L 0.270 41.5 138.4 2.3 3.5 0.244 0.679
BLIP-L+M 0.259 41.9 139.6 2.1 3.1 0.206 0.682
$-r_{f}$ 0.267 43.0 142.3 2.8 4.4 0.249 0.691
$-r_{a}$ 0.257 41.1 132.9 1.5 2.3 0.174 0.66
$-r_{kl}$ 0.241 27.6 98.9 1.4 1.9 0.135 0.62
$-ppo$ 0.287 39.4 127.6 2.5 3.76 0.212 0.664
@@
Refer to caption A man sitting on a chair with a stuffed animal, specifically a teady bear, on his lap a man sitting on a chair holding a large stuffed animal
Figure 13: LLaVa-RLHF vs. MOCHa. We illustrate that RLHF training does not necessarily solve the hallucination problem of VLM models by showing a generation produced by LLaVa-RLHF Sun et al. (2023a) compared to BLIP+MOCHa. For both models, we use the prompt “a photography of" for generation. See Table 10 for a quantitative comparison.
@@ Appendix C Additional Results
C.1 Full Quantitative Results
We show in Table 6 the full results, comparing the MOCHa optimized models (marked by +M) to the baselines (Figure 7 was prepared using this data).
[…]
Dedicated
UD-L+OccXE 33.9 27.0 110.7 5.9 3.8
UD-L+OccSC 37.7 28.7 125.2 5.8 3.7
CIICXE 37.3 28.5 119.0 5.3 3.6
CIICSC 40.2 29.5 133.1 7.7 4.5
COSNetXE 39.1 29.7 127.4 4.7 3.2
COSNetSC 42.0 30.6 141.1 6.8 4.2
End-to-end
BLIP 41.5 31.1 138.4 3.5 2.3
BLIP-2 43.4 31.7 144.3 2.6 1.7
Table 8: Older dedicated methods for reduced-hallucination captioning vs. end-to-end modern VLMs for image captioning. Results are given on the Karpathy test split of MS-COCO dataset, including closed-vocabulary hallucination metrics as commonly reported by such dedicated methods. B@4, C, M, CH denote BLEU-4, CIDEr, METEOR, and CHAIR metrics respectively. We see that older, dedicated methods with weaker backbones are outperformed by modern VLMs on all metrics, including the smaller BLIP(-Large) and the larger BLIP-2(-2.7B). XE and SC indicate cross-entropy and SCST (RL) optimization respectively. Best and second-best metric values are shown in bold and underlined text respectively.
C.2 Comparisons of OpenCHAIR and CHAIR
In Tables 34 we provide full numeric results for our human evaluation of OpenCHAIR and CHAIR across a variety of captioning model predictions, as we discuss in the main paper.
In Figure 14, we illustrate the number of unique object types found in these benchmarks. We note that OpenCHAIR contains a much larger diversity of object types, even when considering the full contents of CHAIR’s synonym list.
Figure 14: Object Type Coverage, CHAIR vs. OpenCHAIR. We display the object type coverage of CHAIR (over MS-COCO) and OpenCHAIR, measured as the number of unique objects. In OPENChair, objects are found using the parsing method described in Section B.2. As can be observed, the proposed benchmark has significantly greater coverage of different objects.
C.3 Additional Ablations
[…]
BLIP 41.5 138.4 2.3 3.5 0.246 0.679
BLIP+M 41.9 139.6 2.1 3.1 0.206 0.682
$-r_{f}$ 43.0 142.3 2.8 4.4 0.249 0.691
$-r_{a}$ 41.1 132.9 1.5 2.3 0.174 0.66
Table 9: Reward Ablation. We ablate the effect of the fidelity $r_{f}$ and adequacy $r_{a}$ terms in our reward function, finding that using each alone significantly degrades performance with respect to hallucinations or textual quality.
Refer to caption Refer to caption Refer to caption
$\emptyset$ This is a picture of a large old fashioned car that was parked by a group of people People at festival standing around in open field
$-r_{f}$ A car parked in the grass with a surfer standing near it A woman standing next to a herd of animals with an umbrella
$-r_{a}$ Spectators could enjoy the old fashions of the fifties That are some very nice people who are very fun to view them
$r$ A vintage car parked on a field next to people A young man with a large umbrella next to a herd of animals
Reward Ablations. In Table 9, we provide numeric results for ablating the fidelity and adequacy terms in our reward function. As discussed in the main paper, removing either of these reward terms leads to a degradation with respect to either hallucinations or textual quality, while using both together displays a synergistic effect with hallucinations reduced (as reflected by metrics such as CHAIR) while preserving or even improving caption quality (as reflected by general textual quality metrics such as BLEU-4). We also show a qualitative illustration of these results in Figure 2.